The 2008 balance sheet of Saddle Creek, Inc., showed current assets of $1,510 and current liabilities of $960. The 2009 balance sheet showed current assets of $1,620 and current liabilities of $1,000. What was the company's 2009 change in net working capital, or NWC? Multiple Choice $-150 $150 $2,030 $-180 $70
Answer #1

Net Working Capital = Current Assets - Current Liabilities

2008 working capital = 1,510-960

=$550

2009 Working Capital = 1,620-1,000=$620

Change in net Working Capital = 2009 working capital - 2008 working capital

= $620-$550

=$70